A 100g object floats in a liquid of relative density 1.2.calculate

a)mass of the liquid displaced by the object.



1
Expert's answer
2021-02-04T11:43:04-0500

By the definition, the buoyant force is equal to the weight of the liquid displace:


"F_B=\rho_{liquid}V_{liquid}g,""m_{liquid}g=\rho_{liquid}V_{liquid}g,""m_{liquid}=\rho_{liquid}V_{liquid}."

We can find "\rho_{liquid}" from the definition of relative density:


"RD=\dfrac{\rho_{liquid}}{\rho_{water}},""\rho_{liquid}=RD\rho_{water}=1.2\cdot1000\ \dfrac{kg}{m^3}=1200\ \dfrac{kg}{m^3}."

Since the floating object displaces the volume of water equivalent to its own volume, we can write:


"m_{liquid}=\rho_{liquid}V_{liquid}=m_{liquid}=\rho_{liquid}V_{obj}."

Unfortunately, we don't know the volume of the object. Let's suppose that the volume of the object equals "10\ cm^3".

Then, we get:


"m_{liquid}=1200\ \dfrac{kg}{m^3}\cdot10\ cm^3\cdot(\dfrac{1\ m}{100\ cm})^3=1.2\cdot10^{-2}\ kg."